# Simultaneous Reward Distillation and Preference Learning: Get You a Language Model Who Can Do Both

### v1

@@ Abstract
Reward modeling of human preferences is one of the cornerstones of building usable generative large language models (LLMs). While traditional RLHF-based alignment methods explicitly maximize the expected rewards from a separate reward model, more recent supervised alignment methods like Direct Preference Optimization (DPO) circumvent this phase to avoid problems including model drift and reward overfitting. Although popular due to its simplicity, DPO and similar direct alignment methods can still lead to degenerate policies, and rely heavily on the Bradley-Terry-based preference formulation to model reward differences between pairs of candidate outputs. This formulation is challenged by non-deterministic or noisy preference labels, for example human scoring of two candidate outputs is of low confidence. In this paper, we introduce DRDO (Direct Reward Distillation and policy-Optimization), a supervised knowledge distillation-based preference alignment method that simultaneously models rewards and preferences to avoid such degeneracy. DRDO directly mimics rewards assigned by an oracle while learning human preferences from a novel preference likelihood formulation. Our experimental results on the Ultrafeedback and TL;DR datasets demonstrate that policies trained using DRDO surpass previous methods such as DPO and e-DPO in terms of expected rewards and are more robust, on average, to noisy preference signals as well as out-of-distribution (OOD) settings.
@@ 1 Introduction
In the development of large language models (LLMs), robust modeling of human preferences is essential for producing outputs that are contextually and situationally appropriate, in order to result in models that users will actually want to use. Inherent in this problem is that while popular approaches like direct preference optimization (DPO) and its variants implicitly assume that pairs of preferred and dispreferred samples in preference data have an unambiguous winner, this does not reflect the reality of actual data, where human-annotated preferences may have low labeler confidence or the preference strength itself might be weak in the data. As such, reward functions estimated on such data with current popular approaches can lead to policy degeneracy and underfitting the true preference distribution. To address these challenges, we follow the insight that certain problems with the DPO implicit reward formulation at the policy learning stage are less problematic at the reward modeling stage, and make the following novel contributions: 1) We introduce Direct Reward Distillation and policy-Optimization (DRDO), a novel approach preference optimization approach that combines the two stages by explicitly distilling rewards into the policy model (Fig. 1); 2) We provide a thorough theoretical grounding of problems with DPO and variants according to the Bradley-Terry model that demonstrates why they are challenged by nuanced or “non-deterministic” preference pairs; 3) Through experiments on the Ultrafeedback and TL;DR datasets along with benchmarks like AlpacaEval, we demonstrate that DRDO outperforms competing popular preference optimization methods across multiple models and sizes at the aggregate data level, and is better able to distinguish preferred and dispreferred responses with low confidence or small reward difference scores.
Figure 1: Unlike popular supervised preference alignment algorithms like Direct Preference Optimization (DPO; Rafailov et al. (2024)) that learns rewards implicitly, DRDO directly optimizes for explicit rewards from an Oracle while simultaneously learning diverse kinds of preference signals during alignment. Optimized with a simple regression loss based on difference of rewards assigned by the Oracle and a novel focal-log-unlikelihood component (see Section 4), DRDO avoids DPO’s particular challenges at learning non-deterministic preference pairs, thereby bridging the gap between estimating the preference distribution from the data and the true preference distribution $p^{*}$. Additionally, DRDO does not require an additional reference model during training and can leverage reward signals even when preference labels are not directly accessible.
@@ 2 Background and Related Work
Recent advancements in Large Language Models (LLMs) involve refining pretrained LLMs for downstream tasks by utilizing human-written completions (Chung et al., 2022; Mishra et al., 2021) or datasets labeled with human-preferred completions and contrasting alternatives (Ouyang et al., 2022; Bai et al., 2022; Ziegler et al., 2019). The two most prominent techniques in learning from preference data are reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO). In the following sections, we summarize these methods.
@@
Reinforcement Learning from Human Feedback (RLHF) aims to harmonize LLMs with human preferences and values (Christiano et al., 2017). Conventional RLHF typically consists of three phases: supervised fine-tuning, reward model training, and policy optimization. Proximal Policy Optimization (PPO; Schulman et al. (2017)) is a widely used algorithm in the third phase of RLHF. RLHF has been extensively applied across various domains, including mitigating toxicity (Korbak et al., 2023; Amini et al., 2024), addressing safety-concerns (Dai et al., 2023), enhancing helpfulness (Tian et al., 2024), web search and navigation (Nakano et al., 2021), and enhancing reasoning in models (Havrilla et al., 2024). Recent work by (Casper et al., 2023) has identified challenges and problems throughout the entire RLHF pipeline, from gathering preference data to model training to biased results, such as verbose outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
@@
where $\sigma(\cdot)$ is the logistic function, and $r(x,y)$ represents an underlying reward function. To estimate this reward function $r$, RLHF minimizes the negative log-likelihood of the data:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log(\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})))\right].$ (2)
@@
Given the intricacy and complexity of online preference optimization (Zheng et al., 2023), research has proliferated into more efficient and simpler offline algorithms. Direct Preference Optimization (DPO) (Rafailov et al., 2024) is a notable example, which demonstrates that the same KL-constrained objective as RLHF can be optimized without explicitly learning a reward function. The problem is reformulated as a maximum likelihood estimation (MLE) over the distribution $\pi_{\theta}$, leading to the following objective:
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\text{ref}}(y_{w}|x)}\right)\right]$ (3)
where $\beta$ is a regularization term controlling the KL-constraint strength. In this case, the implicit reward function is $r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$. Rafailov et al. (2024) further showed that all reward models based on the Plackett-Luce distribution (Plackett, 1975; Luce, 2005), including Bradley-Terry, can be expressed in this form. However the absence of an explicit reward model in DPO limits its ability to sample preference pairs from the optimal policy. Zhao et al. (2023) have investigated augmenting preference data using a trained SFT policy and Liu et al. (2024b) have done the same with a refined SFT policy with rejection sampling, enabling policy learning from data generated by the optimal policy.
Exploring Preference Optimization Objectives
Various other preference optimization objectives have been proposed in the literature. Ranking objectives facilitate compare multiple instances beyond just pairs (Dong et al., 2023; Liu et al., 2024a; Song et al., 2024; Yuan et al., 2023). Hong et al. (2024) and Xu et al. (2023) investigate objectives that do not require a reference model. Bansal et al. (2024) suggest a method that jointly optimizes instructions and responses, demonstrating notable improvements over DPO. Zheng et al. (2024) focuses on post-training extrapolation between the (SFT) model and the aligned model to enhance overall performance. In contrast to MLE-based DPO, Fisch et al. (2024) argue, given potentially unlimited unlabeled samples $(x,y_{1},y_{2})\sim\rho$, a simple “distillation” loss that explicitly regresses the implicit reward difference in DPO onto the true pointwise reward estimates provided by an explicit reward model or a suite of reward models [^1] Notably Fisch et al. (2024) propose two versions of this reward distillation loss, one with a single reward model, the other with an ensemble of reward models. For simplicity of analysis, we only show the former here. For all our experiments with e-DPO baseline, we use the ensembled version of this loss., as follows:
$\mathcal{L}_{\mathrm{distill}}(r^{*},\pi_{\theta};\rho)=\mathbb{E}_{\rho(x,y_{1},y_{2})}\left[\left(r^{*}(x,y_{1})-r^{*}(x,y_{2})-\beta\log\frac{\pi_{\theta}(y_{1}\mid x)\pi_{\text{ref}}(y_{2}\mid x)}{\pi_{\theta}(y_{2}\mid x)\pi_{\text{ref}}(y_{1}\mid x)}\right)^{2}\right].$ (4)
Iterative sampling algorithms like rejection sampling are compute-intensive, and in this work, we focus exclusively on offline settings, avoiding any iterative training processes. For simplicity, we define $\pi_{\text{ratio}}(y|x)=\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ which can be parametrized with the winning and losing responses to get the respective likelihood ratios. We build upon insights from the previous literature to bypass issues inherent in existing offline preference optimization frameworks.
3 Theoretical Motivation: Limitations of Direct Alignment
In order to theoretically motivate DRDO, we first provide an in-depth analysis of limitations of DPO specific to non-deterministic or weak preference labels, as well as limitations of recent proposed solutions like IPO (Azar et al., 2023) and Ensemble-DPO (e-DPO; Fisch et al. (2024)). We then apply knowledge distillation-based techniques on rewards to provide a practical algorithm for DRDO.
Non-Deterministic Human Preferences
Following Bradley & Terry (1952), Rafailov et al. (2024) and other preference optimization frameworks posit that the relative preference of one outcome over another is governed by the true reward differences, expressed as $p^{*}(y_{1}\succ y_{2})=\sigma(r_{1}-r_{2})$, where $p^{*}$ is the true preference distribution. Generally, in the RLHF framework, the true preference distribution is typically inferred from a dataset of human preferences, using a reward model $r$ that subsequently guides the optimal policy learning. More importantly, to estimate the rewards and thereby the optimal policy parameters, the critical reward modeling stage involves human annotators choosing between pairs of candidate answers $(y_{1},y_{2})$, indicating their preferences. [^2] This framework can be extended to rank multiple responses using the Plackett-Luce (Plackett, 1975) model. As such, typical alignment methods assume that $p(y_{w}\succ y_{l}|x)$ (the human annotations of preference) is equivalent to $p^{*}(y_{1}\succ y_{2}|x)$ or any ranking or choice thereby established with the human decisions. However, prospect theory and empirical studies in rational choice theory suggest that human preferences are often stochastic, intransitive, and can fluctuate across time and contexts (Tversky, 1969; von Weizsäcker, 2005; Regenwetter et al., 2011).
Existing direct alignment methods, such as DPO-based supervised alignment, assume access to deterministic preference labels, disregarding the inherent variability in human judgments, even when popular preference datasets are inherently annotated with such variability, noise, or “non-deterministic preferences” given their provenance in human labeling. More importantly, such implicit trust in the preference data by DPO-like algorithms without explicit instance-level penalization on the loss, can cause policies that are trained to deviate from true intentions of human preference learning (see Lemma 1 and Lemma 2 for details). Additionally, in many datasets, a significant proportion of preference pair annotations display low human confidence, or are scored the same by common reward models (e.g., GPT-4) despite being textually different, indicating that the two responses are likely semantically similar. As with policy learning, such discrepancies in preference signals can similarly derail reward learning and limit reward models from reaching a consensus, even with majority voting with reward ensembles (Wang et al., 2024a). These cases reflect the stochastic nature of human choices, and challenge the assumption of stable, deterministic preferences in alignment frameworks.
Assumption 1.
Let $\mathcal{D}_{\text{pref}}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$ be an offline dataset of pairwise preferences with sufficient coverage, where each $x^{(i)}$ is a prompt, and $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are the corresponding preferred and dispreferred responses, respectively. Let $r^{*}(x,y)\in\mathbb{R}$ be an underlying true reward function that is deterministic333 Deterministic and non-deterministic preferences are only defined on the true preference distribution $p^{*}$ and should not be confused with the empirical probabilities or confidence assigned by the policy. The use of ”deterministic” here is simply to imply that the true reward function $r^{*}(x,y)$ is finite and scalar. and finite everywhere. Let $\pi_{\theta}^{*}(y\mid x)$ be the learned model and $\pi_{\text{ref}}(y\mid x)$ the reference, with $\text{supp}(\pi_{\text{ref}})=\mathcal{Y}$. Assume $\text{supp}(\rho)=\text{supp}(\mu)\times\mathcal{Y}\times\mathcal{Y}$, where $\mathcal{Y}$ is the space of all responses, $\rho$ is the data distribution, and $\mu$ is the prompt or context distribution.
Proposition 1 (Non-Deterministic Preferences).
Define the subset $\mathcal{D}_{\text{nd}}\subset\mathcal{D}_{\text{pref}}$ with non-deterministic preferences as $\mathcal{D}_{\text{nd}}=\{(x,y,y^{\prime})\in\mathcal{D}_{\text{pref}}\mid P(y\succ y^{\prime}\mid x)\approx\frac{1}{2}\}$, where $|\mathcal{D}_{\text{nd}}|\ll N$. Under Assumption 1 and antisymmetric preferences (Munos et al., 2023), the preference relation is given by $P(y_{w}^{(i)}\succ y_{l}^{(i)}\mid x^{(i)})=1-P(y_{l}^{(i)}\succ y_{w}^{(i)}\mid x^{(i)})$. Given non-deterministic preferences, i.e., $(y,y^{\prime})\in\mathcal{D}_{\text{nd}}$, the Bradley-Terry model assigns $\Delta r=0$, implying $r^{*}(x,y)=r^{*}(x,y^{\prime})$. Thus, $\forall(x,y,y^{\prime})\in\mathcal{D}_{\text{nd}},\ \Delta r=0$. See Section A.1 for a complete derivation.
Lemma 1.
Under Proposition 1, a) the DPO implicit reward difference in its objective $\frac{\pi_{\theta^{*}}(y)\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\pi_{\mathrm{ref}}(y)}\to 0$, that leads to the policy empirically underfitting the preference distribution. b) For $|\mathcal{D}_{\text{nd}}|\ll N$ where $N$ is finite, if DPO estimates that $p^{*}(y\succ y^{\prime})=1$, then $\frac{\pi_{\theta^{*}}(y)\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\pi_{\mathrm{ref}}(y)}\to\infty$. c) For all minimizers $\pi_{\theta^{*}}$ of the DPO objective (Eq. 3), it follows that $\pi_{\theta^{*}}(y_{l})\to 0$ and $\pi_{\theta^{*}}(\mathcal{C}(y_{l})^{c})\to 1$, where $\mathcal{C}(y_{l})^{c}$ denotes the complement of the set of dispreferred responses $y_{l}^{(i)}$, $\forall i\in\mathbb{N}$.
Given non-trivial occurrences of non-deterministic preference pairs in typical preference learning datasets, a consequence of Lemma 1 is that DPO’s learned optimal policy can effectively assign non-zero or even very high probabilities to tokens that never appear as preferred in the training data, causing substantial policy degeneracy. Moreover, as noted and shown emprirically in previous work (Azar et al., 2023; Pal et al., 2024), DPO effectively underfits the preference distribution because its empirical preference probabilities (RHS of Eq. 3 without the expectation) are only estimates of the true preference probabilities, especially when $p^{*}(y\succ y^{\prime})\in\{0,1\}$. A noteworthy implication of Lemma 1 is that this weak regularization effect of DPO can theoretically assign very high probabilities to the complement set of dispreferred tokens that never appear in the training data at all, especially when $|N_{\text{nd}}|\ll N$ for finite data regimes. In realistic settings where non-deterministic preferences constitute a non-trivial proportion of data, Lemma 1 additionally implies that DPO leads to unstable updates and inconsistent policy behavior, where the gradient update is effectively cancelled out for these samples since the log probabilities of both the winning and the losing responses are roughly equal ($\Delta r\approx 0$), so the scaled weighting factor (sigmoid of implicit reward differences) does not contribute as much as when $p^{*}(y\succ y^{\prime})\in(0,1)$. In other words, with DPO, $\pi_{\theta^{*}}$ not only sees less of this type of preference but also fails to adequately regularize when it does.
However, these approaches have inherent limitations. IPO regresses the log-likelihood difference on a Bernoulli-distributed preference label, failing to capture nuanced strength in relative preferences. Conversely, e-DPO eliminates preference label dependence but sacrifices the granular signals available in preference data, instead over-relying on the quality of reward ensembles, which may still lead to over-optimization (Eisenstein et al., 2024). Furthermore, the use of reward ensembles in e-DPO introduces significant computational overhead, potentially limiting its broader applicability due to increased resource requirements. Consider the following lemma that derives from Assumption 1 and Proposition 1:
Lemma 2.
Under Proposition 1 and in the spirit of Fisch et al. (2024)’s argument, using e-DPO alignment over non-deterministic preferences pairs leads to $\frac{\pi_{\theta^{*}}(y)\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\pi_{\mathrm{ref}}(y)}\to\infty$ for $(y,y^{\prime})\in\mathcal{D}_{nd}$ where $y=y_{w}^{(i)}$ and $y^{\prime}=y_{l}^{(i)}$, $\forall i\in\mathbb{N}$. Then, for all minimizers $\pi_{\theta^{*}}$ of the e-DPO objective (Eq. 4), it follows that $\pi_{\theta^{*}}(\mathcal{C}(y_{l})^{c})\to 1$ with $0<\pi_{\theta^{*}}(y_{w}^{(i)})\leq 1$, $\forall i\in\mathbb{N}$, where $\mathcal{C}(y_{l})^{c}$ denotes the complement of the set of responses $y$ appearing as preferred $y_{l}^{(i)}$, $\forall i\in\mathbb{N}$.

### v2

@@ Abstract
Traditional RLHF-based LLM alignment methods explicitly maximize the expected rewards from a separate reward model. More recent supervised alignment methods like Direct Preference Optimization (DPO) circumvent this phase to avoid problems including model drift and reward overfitting. Although popular due to its simplicity, DPO and similar direct alignment methods which rely heavily on the Bradley-Terry-based pairwise preference formulation can still lead to degenerate policies when challenged by non-deterministic or noisy preference labels, for example human scoring of two candidate outputs with low confidence. This paper introduces DRDO (Direct Reward Distillation and policy-Optimization), which simultaneously models rewards and preferences to avoid such degeneracy. DRDO directly mimics rewards assigned by an oracle while learning human preferences with a novel preference likelihood formulation. Results on the Ultrafeedback and TL;DR datasets demonstrate that DRDO-trained policies surpass methods such as DPO and e-DPO in terms of expected rewards and are more robust, on average, to noisy preference signals as well as out-of-distribution (OOD) settings.
Machine Learning, ICML
@@ 1 Introduction
In the development of large language models (LLMs), robust modeling of human preferences is essential for producing models that users actually find useful. Inherent in this problem is that while popular approaches like Direct Preference Optimization (DPO) and its variants implicitly assume that pairs of preferred and dispreferred samples in preference data have an unambiguous winner, this does not reflect the reality of actual data, where human-annotated preferences may have low labeler confidence or the preference strength itself might be weak. As such, reward functions estimated on such data with current popular approaches can lead to policy degeneracy and underfitting the true preference distribution. To address these challenges, we follow the insight that certain problems with the DPO implicit reward formulation at the policy learning stage are less problematic at the reward modeling stage, and make the following novel contributions: 1) We introduce Direct Reward Distillation and policy-Optimization (DRDO), a novel efficient, non-ensemble, reference-free method for preference optimization that combines the two stages by explicitly distilling rewards into the policy model (Fig. 1); 2) We provide a theoretical and practical grounding of problems with DPO and variants that assume the Bradley-Terry model, demonstrating why they are challenged by nuanced or “non-deterministic” preference pairs, and how DRDO avoids similar limitations; 3) Through experiments on the Ultrafeedback and TL;DR datasets and AlpacaEval benchmark, we demonstrate that DRDO outperforms competing popular preference optimization methods across multiple models and sizes at the aggregate data level, and at the sample level is better able to distinguish preferred and dispreferred responses even with low annotator confidence or small reward differences.
Figure 1: Unlike popular supervised preference alignment algorithms like Direct Preference Optimization (DPO; Rafailov et al. (2024b)) that learns rewards implicitly, DRDO directly optimizes for explicit rewards from an Oracle while simultaneously learning diverse kinds of preference signals during alignment. Optimized with a simple regression loss based on difference of rewards assigned by the Oracle and the introduction of a focal-log-unlikelihood component (see Sec. 4), DRDO avoids DPO’s particular challenges at learning non-deterministic preference pairs, thereby bridging the gap between the preference distribution estimated from the data and the true preference distribution $p^{*}$. Additionally, DRDO does not require an additional reference model during training and can leverage reward signals even when preference labels are not directly accessible.
@@ 2 Background and Related Work
Recent advancements in LLMs involve refining pretrained models for downstream tasks by utilizing human-written completions (Chung et al., 2022; Mishra et al., 2021) or datasets labeled with human-preferred completions and contrasting alternatives (Ouyang et al., 2022; Bai et al., 2022; Ziegler et al., 2019). The two most prominent techniques in learning from preference data are Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO). We summarize these methods below.
@@
Reinforcement Learning from Human Feedback (RLHF) aims to harmonize LLMs with human preferences and values (Christiano et al., 2017). Conventional RLHF typically consists of three phases: supervised fine-tuning, reward model training, and policy optimization. Proximal Policy Optimization (PPO; Schulman et al. (2017)) is a widely used algorithm in the third phase of RLHF. RLHF has been extensively applied across various domains, including mitigating toxicity (Korbak et al., 2023; Amini et al., 2024), addressing safety-concerns (Dai et al., 2023), enhancing helpfulness (Tian et al., 2024), web search and navigation (Nakano et al., 2021), and enhancing reasoning in models (Havrilla et al., 2024). Casper et al. (2023) identified challenges and problems throughout the entire RLHF pipeline, from gathering preference data to model training to biased results such as verbose outputs (Dubois et al., 2024; Singhal et al., 2023; Wang et al., 2023).
@@
where $\sigma(\cdot)$ is the logistic function, and $r(x,y)$ the underlying reward function. To estimate this reward function $r$, RLHF minimizes the negative log-likelihood of the data:
$\mathcal{L}_{R}(r_{\phi},\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\\ \left[\log(\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})))\right].$ (2)
@@
Given the intricacy and complexity of online preference optimization (Zheng et al., 2023), research has proliferated into more efficient and simpler offline algorithms. Direct Preference Optimization (DPO; Rafailov et al. (2024b)) is a notable example, which demonstrates that the same KL-constrained objective as RLHF can be optimized without explicitly learning a reward function. The problem is reformulated as a maximum likelihood estimation (MLE) over the distribution $\pi_{\theta}$, leading to the following objective:
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\\ \left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\text{ref}}(y_{w}|x)}\right)\right]$ (3)
where $\beta$ is a regularization term controlling the KL-constraint strength and $r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ is the implicit reward function. With the growing focus on offline alignment, recent work has examined preference underfitting in this setting, proposing various solutions. These include learning from a confidence set of rewards (Fisch et al., 2024), adopting a general preference model (Munos et al., 2023; Azar et al., 2024), or applying a straightforward regularization of the original DPO objective (Pal et al., 2024a). Notably, all these approaches rely on a reference model, not only for expressing the optimal policy as the analytical solution to the minimum relative entropy problem (Ziebart et al., 2008; Peng et al., 2019) but also to stabilize training by constraining the policy distribution to remain close to the reference model. While theoretically sound, Munos et al. (2023) suggests this constraint can make policies prioritize high rewards over truly learning human preferences.
As such, certain works avoid this constraint and focus on lightweight “reference model-free” solutions using a careful construction of DPO-inspired Bradley-Terry (BT)-based implicit rewards—such as logit-based (Hong et al., 2024), length-normalized (Meng et al., 2024) as well as un-normalized implicit rewards (Xu et al., 2024)—while still showing high performance in preference alignment benchmarks including cross-task settings like machine translation. Similarly, efforts to learn policies from general preference models aim to reduce dependence on the sampling distribution, a key limitation of Bradley-Terry models. Munos et al. (2023), Rosset et al. (2024), and Calandriello et al. (2024) use game-theoretic “online” approaches for robustness, while Choi et al. (2024) propose a Chain-of-Thought (CoT) policy with pairwise conditioning to improve alignment. However, these methods often suffer from sample inefficiency due to iterative sampling from approximate geometric mixtures of policies (Rosset et al., 2024). While these approaches mitigate dependence on the sampling distribution, they do not model non-deterministic preferences in policy training. In contrast, DRDO adaptively learns such preferences while decoupling its learning from a reference model-based KL-divergence constraint.
Preference Optimization Objectives
Several preference optimization objectives have been explored. Ranking objectives extend comparisons beyond pairs (Dong et al., 2023; Liu et al., 2024; Song et al., 2024; Yuan et al., 2023), while Hong et al. (2024); Xu et al. (2023) propose reference-free methods. Bansal et al. (2024) optimize instructions and responses jointly, improving on DPO, and Zheng et al. (2024) enhances post-training extrapolation between SFT and aligned models. By comparison, DRDO derives its objective from knowledge distillation (Hinton, 2015) and focal-loss literature (Lin et al., 2018; Yi et al., 2020) with a novel contrastive log-unlikelihood objective that learns preferences adaptively.
3 Motivation for DRDO
In order to motivate DRDO, we first discuss theoretical and practical limitations of DPO specific to non-deterministic or weak preference labels. For space reasons, proofs are deferred to the Appendix.
Problem Formulation
Let $\mathcal{D}_{\text{pref}}=\{(x,y_{w},y_{l})\}_{i=1}^{N}$ be an offline dataset of pairwise preferences with sufficient coverage, where $y_{w}$ and $y_{l}$ are the respective winning (preferred) and losing (less preferred) completions given a context $x$. In contrast to deterministic preferences, which are defined as those where $p^{*}\in\{0,1\}$ (Fisch et al., 2024), let $\mathcal{D}_{\text{nd}}\subset\mathcal{D}_{\text{pref}}$ denote the subset of non-deterministic preference pairs where $P(y_{w}\succ y_{l}|x)\approx\frac{1}{2}$. These cannot be perfectly captured by Bradley-Terry models but are prevalent in popular preference alignment datasets. Assume three possible completions $y_{1},y_{2},y_{3}\in\mathcal{Y}$ where $\mathcal{Y}$ is the space of all possible completions. Let $r^{*}(x,y)\in\mathbb{R}$ be an underlying true reward function that is deterministic and finite for all completions, $\pi_{\theta^{*}}(y|x)$ be the learned policy, and $\pi_{\text{ref}}(y|x)$ be the reference policy with $\text{supp}(\pi_{\text{ref}})=\mathcal{Y}$. Given $\text{supp}(\rho)=\text{supp}(\mu)\times\mathcal{Y}\times\mathcal{Y}$ where $\rho$ is the data distribution and $\mu$ is the context distribution, and given potentially biased sampling of preference pairs from $\mathcal{Y}$ to construct $\mathcal{D}_{\text{pref}}$, the problem is to learn a policy $\pi_{\theta^{*}}$ that effectively handles both deterministic and non-deterministic preferences in an offline fashion.
The key motivation behind DRDO is to address non-deterministic preferences while maintaining performance across general preference data. Existing supervised “offline” methods such as DPO or e-DPO assume preferences can be specified with a Bradley-Terry model and that policies can be optimized by inferring the Bradley-Terry model directly from preference data. At the core, these algorithms refine the Supervised Fine-tuned (SFT or reference) model from which the trainable policy $\pi_{\theta}$ is typically initialized, and they both apply a soft-constraint in this optimization where $\pi_{\theta}$ is forced to remain distributionally close to $\pi_{\text{ref}}$ via a KL-penalty term.
This soft-constraint—while well-motivated for training stability and to reduce $\pi_{\theta}$’s likelihood of assigning high-probability to actions not supported by $\pi_{\text{ref}}$ (Ziebart et al., 2008; Peng et al., 2019)—creates two specific problems. Firstly, it leads to divergent objectives during preference alignment. We will call this the misalignment problem. Secondly, practical alignment methods like DPO and e-DPO apply this soft-constraint with a regularization-strength parameter $\beta$ that offers only partial support in preventing DPO-aligned policies from inducing degenerate solutions, especially for deterministic [^1] Prior work (Azar et al., 2024; Fisch et al., 2024; Pal et al., 2024b) shows that DPO policy degeneracy arises in finite-sample settings with deterministic preferences, independent of $\beta$. preferences. In this work, we show how these limitations can be equally applicable in the case of non-deterministic preferences.
Misalignment of Rewards and Preferences
The misalignment problem is where the reward-optimal policy, or the policy that maximizes Elo-score under the BT assumption, is fundamentally divergent from the preference optimal policy, or the policy that maximizes the probability of selecting the winning response (Munos et al., 2023). In other words, traditional reward modeling assumes that maximizing reward leads to optimal behavior, but in preference-based learning, the best winning policy can be different from the highest-reward policy. We argue that in supervised learning algorithms like DPO that combine these stages, this divergence is amplified in the presence of non-deterministic preferences in training data since such preferences cannot be perfectly captured with a BT model, leading to reward misspecification (full derivation in Section B.1).
This misalignment also corresponds to the induced BT model’s explicit dependence on the sampling distribution or the behavior policy (Munos et al., 2023) (see Proposition 1 in Section B.1 for a full exposition). Under sampling biases in preference data collection, the optimal policy’s reward misspecification can be highly localized. Consider a case where preferences between completions $y_{2}$ and $y_{3}$ are non-deterministic (i.e., $P(y_{2}\succ y_{3})=P(y_{3}\succ y_{2})=\frac{1}{2}$). While an ideal Bradley-Terry model should assign equal rewards to such completions, the reality of data collection often introduces skewed sampling distributions. For instance, if the training data contains predominantly $y_{1}$-$y_{2}$ comparisons (Choi et al., 2024), the model may accurately estimate rewards for these frequently compared completions while misspecifying only $y_{3}$’s reward. Such localized misspecification presents distinct challenges: in off-policy training, DPO implicitly trusts the preference dataset as ground truth, while in on-policy settings, non-deterministic preferences are difficult for non-expert annotators to identify (Stiennon et al., 2020), making it challenging to detect and filter problematic completions in both scenarios. This challenge is magnified in modern LLM-assisted annotation pipelines; for instance, our examination of the Ultrafeedback datasets (Cui et al., 2024) reveals that in approximately 17% of preference pairs, the candidate responses have equal scores.
The core issue stems from the inherent tension in scalable oversight (Amodei et al., 2016): while it successfully increases the volume of preference signals, it provides no clear-cut mechanism to control the underlying sampling distribution that generates these pairs. This fundamentally limits approaches that rely on inferring rewards through Bradley-Terry models. Previous work (Choi et al., 2024) attempts to resolve this misalignment using an optimal Chain-of-thought (CoT) policy with additional conditioning on pairwise interactions that does not depend on the sampling distribution, while others (Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024) use game-theoretic “online” approaches that learn from a general preference model that is robust to sampling, instead of a pointwise reward model. However, these approaches require iterative sampling from approximate geometric mixtures of policies (Rosset et al., 2024). Additionally, all the aforementioned require a reference model, which still does not resolve misalignment created by the KL-penalty soft constraint.
Practical Limitations in Learning Non-deterministic Preferences
The misalignment issue illustrated above applies to any induced BT model, but algorithms like DPO present separate practical challenges. In particular, DPO estimates the preference distribution using an implicit reward function $r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ that is effectively a function of $\pi_{\theta}$ where $\pi_{\text{ref}}$ is not updated during training but provides a soft-KL constraint, weighted by $\beta$. For two completions $y$ and $y^{\prime}$ with a non-deterministic preference relation, one can see that when DPO estimates $p^{*}(y\succ y^{\prime})$ using $\sigma\big{(}\beta\log\frac{\pi_{\theta}(y|x)\pi_{\text{ref}}(y^{\prime}|x)}{\pi_{\theta}(y^{\prime}|x)\pi_{\text{ref}}(y|x)}\big{)}$, this leads $\pi_{\theta}$ to assign equal rewards to both, meaning it learns that $r^{*}(x,y)=r^{*}(x,y^{\prime})$ for any $(x,y,y^{\prime})\in\mathcal{D}_{\text{nd}}$. This implies that the reward model fails to differentiate between completions in these cases, especially when $\pi_{\text{ref}}$ is not fully optimized or is uniform ($\pi_{\text{ref}}\sim\mathcal{U}(\mathcal{Y})$(Xu et al., 2024). In other words, with DPO, $\pi_{\theta^{*}}$ not only sees less of this type of preference but also fails to adequately regularize when it does.
A critical implication of the above phenomenon is that DPO’s underfitting persists regardless of the KL-regularization strength, causing the policy to ignore preference signals for non-deterministic pairs (see Section B.3). More concerningly, DPO can assign disproportionately high probabilities to completions that never appear as preferred in training data, particularly when non-deterministic preferences are rare ($|\mathcal{D}_{\text{nd}}|\ll N$). See Section B.4. This is counterintuitive since it reduces likelihood of the preferred completion, $y\sim y_{w}$—which runs counter to the core motivation for preference alignment. More importantly, this policy degeneracy is especially pronounced in large action spaces $\mathcal{Y}$, such as those in LLMs, where the probability boost may not be allocated to actual preferred responses but rather to previously unseen tokens or completions that were never explicitly trained on (Pal et al., 2024a; Rafailov et al., 2024a)—with no clear-cut strategies to recover such probability mass except ad hoc methods like early stopping or a lower learning rate (Fisch et al., 2024).Recent approaches like e-DPO (Fisch et al., 2024) attempt to resolve these issues by casting the DPO objective into a regression problem where the explicit rewards from an external model are directly distilled to DPO’s implicit rewards. However, e-DPO requires careful construction of confidence-sets of reward model ensembles that limits practical usage.
Our above analysis of the problem of misaligned preferences and practical limitations in current methods directly motivate DRDO—we propose a simple offline alignment algorithm that optimizes for rewards and preferences simultaneously. Crucially, DRDO circumvents the misalignment problem by avoiding a reference model-based soft-constraint and casts the reward learning into a distillation problem by explicitly distilling rewards from a preference-converged Oracle. This helps it overcome policy degeneracy issues. As such, by bridging the gap between reward learning and policy optimization, DRDO offers a more principled solution to the preference alignment problem.